Compute the present value of the following: (Round time value factors to 6 decimal places and final answer to the nearest dollar amount. Omit the "$" sign and commas in your response. For example, $123,456 should be entered as 123456.)

Years: 15

Interest Rate: 9.2 percent

Future Value: $607845
Answer #1

Formula: PV = |FV| / (1+R%)^N

PV = 607845 / (1+0.092)^15 = 162350

OR

Using financial calculator BA II Plus - Input details:

#

I/Y = Rate or yield / frequency =

                        9.200000

PMT = Payment = Payment / frequency =

$0.00

N = Total number of periods = Number of years x frequency =

15.00

FV = Future Value =

-$607,845.00

CPT > PV = Present Value of the instrument =

162350